• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name kelleher

    Log entry time 08:04:26 on May 9,2011

    Entry number 358701

    keyword=owl shift summary

    Monday, May 9, 2011

    00:00 Begin shift in controlled access to address target vacuum (entry #358672)
    01:00 Target vacuum much improved (see entry #358676), watching vacuum return to a low enough level to open the gate valves.
    01:30 Asked MCC to open isolation valves on scattering chamber. Vacuum remains steady at 2.5e-5
    02:30 Alexandre starts his low-current tests
    04:30 Access to change gain on BCMs
    06:45 Access to change the amplifier
    07:45 Beam away for beam studies